# MLOmics: Benchmark for Machine Learning on Cancer Multi-Omics Data

### v2

@@ Acknowledgements
center
@@ A.1 MLOmics Structure
In Main Datasets, the repository hosts a comprehensive collection of tasks-ready cancer multi-omics datasets, stored primarily as CSV (Comma-Separated Values) files.
In Baseline and Metrics, the repository provides source codes of baseline models and evaluation metrics for different tasks, typically implemented in Python or R code.
In Downstream Analysis Tools and Resources Linking, the repository encompasses additional tools and resources that complement the main datasets and downstream omics analysis needs, implemented as CSV, PY or R files.
@@
A.2 Dataset Format
MLOmics uses CSV files to manage and store all omics datasets, widely favored in biomedical research, including multi-omics studies. CSV files are plain-text files where data is separated by commas. They maintain a straightforward structure, with rows representing individual data records and columns representing different attributes or variables. Despite its simplicity, CSV remains efficient even with large datasets encountered in genomic and proteomic studies.
For example, consider a simplified CSV file containing mRNA data for a set of gene features (rows) across several patient samples (columns):
Feature Sample1 Sample2 Sample3 Sample4
GeneA 0.23 0.18 0.35 0.21
GeneB 0.56 0.49 0.52 0.58
GeneC 0.19 0.22 0.15 0.17
GeneD 0.08 0.10 0.09 0.12
In this example:
  • Each row corresponds to a specific gene (GeneA, GeneB, GeneC, GeneD).
  • Each column represents a different sample (Sample1, Sample2, Sample3, Sample4).
  • The numeric values in the cells denote the expression levels of each gene in each sample.
Both Python and R provide built-in functions and libraries to read, write, and analyze CSV files efficiently.
A.3 Recruited Cancer
@@ Appendix C MLOmics Feature Scale Processing
To address this, after the typical omics data preprocessing, MLOmics provides three versions of feature scales (Original, Top, and Aligned) to support different machine learning tasks.

### v3

@@ Acknowledgements
Figure 1: Schematic workflow of creating the MLOmics. The process starts with collecting patient samples covering 32 cancer types from the TCGA project. All resources in diverse data types and sizes are uniformly integrated and processed to contain data of four omics types. Datasets for benchmark ML tasks were constructed based on the processed data. MLOmics also selected baselines, metrics, and resources to support downstream biological analysis. Overview of the MLOmics. MLOmics provides an interface for developing and evaluating machine learning models based on cancer multi-omics data. MLOmics provides datasets in three feature scales for 20 classification, clustering, and omics imputation learning tasks. MLOmics also provides statistical, ML, and DL baselines for each task, which are evaluated by fair metrics. Bio-knowledge database linking with MLOmics. MLOmics provides resources to link with other bio-knowledge databases, enabling the integration of outer resources for applications such as ML evaluation, gene-disease association exploration, network inference, and functional analysis.
Figure 2: Schematic of MLOmics resources structure. The MLOmics framework consists of three major components. The Main Datasets file includes all cancer multi-omics datasets for various tasks. The Baseline and Metrics file provides the source code for baseline models and performance metrics. The Downstream Analysis Tools and Resources Linking file offers sources for further analysis and links to additional biological resources.
Figure 3: Experimental results and downstream analyses of machine learning baselines applied to MLOmics datasets. (a) PREC bar plots for each baseline method across all datasets. Overall, machine learning-based methods (Subtype-GAN, DCAP, MAUI, XOmiVAE, CustOmics, and DeepCC) outperformed traditional statistical methods (SVM, XGBoost, RF, and LR). (b) SIL heatmaps for each baseline method across all datasets. Methods employing deep generative neural network architectures (Subtype-GAN, DCAP, MAUI, XOmiVAE, and MCluster-VAEs) generally outperformed other methods (SNF, NEMO, CIMLR, iClusterBayes, and moCluster). (c) Box plots for each baseline method across three imputation datasets. Matrix decomposition methods (SVD, Spectral) outperformed deep learning-based methods (GAIN, GNN). (d–e) Schematic illustrations of downstream analysis results based on the clustering outcomes of XOmiVAE applied to specific cancer patient clustering datasets. In the survival analysis plot, survival curves in different colors correspond to distinct clustering groups. In the volcano plot and KEGG pathway analysis, red and blue indicate downregulated and upregulated genes between patient groups, respectively. In the patient group clustering plot, different colors represent samples belonging to different clusters. In the simulated gene knockout analysis, red and green dots indicate sample clustering before and after expression knockout, respectively.
Appendix
@@ A.1 MLOmics Structure
A.2 Recruited Cancer
@@ Appendix C MLOmics Feature Scale Processing
To address this, after the typical omics data preprocessing, MLOmics provides three versions of feature scales (Original, Top, and Aligned) to support different machine learning tasks. Table 2 summarizes the different feature scales’ details.
@@ Appendix D MLOmics Tasks
The MLOmics database contains multi-omics data curated for a range of task categories, as summarized in Table 3. Definitions for each task category are provided below.
Table 3: Summary of all tasks in the MLOmics under each task category.
Entry Task Categories Task Names
1 Pan-cancer Classification Pan-cancer
2 Golden-standard Subtype Classification GS-BRCA, GS-COAD, GS-GBM, GS-LGG, GS-OV
3 Cancer Subtype Clustering ACC, KIRP, KIRC, LIHC, LUAD, LUSC, PRAD, THCA, THYM
4 Omics Data Imputation Imp-BRCA, Imp-COAD, Imp-GBM, Imp-LGG, Imp-OV